I do not need 66-69!! Anything after that is appreciated!! I promise to rate! Thank you!

Part D: Transition to Parenthood When a married couple has their first baby, is the new child likely to be beneficial or detrimental to the overall quality of their relationship? Feeney, Alexander, Noller, and Hohaus (2003) conducted a study investigating ways in which marriages change during the transition to parenthood. These researchers followed a sample 76 couples over the course of approximately one year. At the beginning of the study, all couples were expecting their first child. Participating couples completed one assessment during the second trimester of pregnancy (time 1), and they completed another assessment six months after the birth of their baby (time 2). As part of each assessment, each spouse completed a measure of marital satisfaction. A paired samples -test can be used to determine if couples marital satisfaction at time 2 (after the birth of the baby) is lower then their satisfaction at time 1 (before the birth of the baby) Results from paired samples t-tests for wives and husbands are reported below Paired Samples Statistics Std. Error Results for wives satisfaction Mean Std. Deiation 4.50 5.15 Mean Pair w sat t1 31.77 29.73 76 76 .52 .59 w sat t2 t1time1 t2 time 2 Paired Samples Correlations Correlation 60 Si Pair 1 w_sat_t1 & w sat_t2 76
Answer #1

66) Mean difference= 31.77-29.73= 2.04

67) standard error of difference= 4.33/sqrt(76)= 4.33/8.72 =0.50

68) The following formula is used to compute the pooled variance:

(n 1)s(n2 1)s n1n2 2 |

1) (4.5)(76 1) (5.15)23.3863 (76 76 76 2

69) Degrees of freedom= n-1= 76-1=75

NOTE: As per Q&A guidelines I have done the first four please re post the rest. Thank you.